# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a variant of reinforcement learning (RL) that learns from human feedback instead of relying on an engineered reward function. Building on prior work on the related setting of preference-based reinforcement learning (PbRL), it stands at the intersection of artificial intelligence and human-computer interaction. This positioning offers a promising avenue to enhance the performance and adaptability of intelligent systems while also improving the alignment of their objectives with human values. The training of large language models (LLMs) has impressively demonstrated this potential in recent years, where RLHF played a decisive role in targeting the model’s capabilities toward human objectives. This article provides a comprehensive overview of the fundamentals of RLHF, exploring the intricate dynamics between machine agents and human input. While recent focus has been on RLHF for LLMs, our survey adopts a broader perspective, examining the diverse applications and wide-ranging impact of the technique. We delve into the core principles that underpin RLHF, shedding light on the symbiotic relationship between algorithms and human feedback, and discuss the main research trends in the field. By synthesizing the current landscape of RLHF research, this article aims to provide researchers as well as practitioners with a comprehensive understanding of this rapidly growing field of research.
Keywords Preference-Based Reinforcement Learning $\cdot$ Human Feedback $\cdot$ Preference Learning $\cdot$ Interactive Learning $\cdot$ Reward Learning $\cdot$ Reward Modeling $\cdot$ Agent Alignment $\cdot$ Human-in-the-Loop $\cdot$ Human-Computer Interaction $\cdot$ Human-Compatible AI
@@ 1 Introduction
In reinforcement learning (RL), an agent traditionally navigates through an environment and attempts to make optimal actions or decisions through a process of trial and error. Whether a decision is optimal or not is determined solely by reward signals. These signals have to be defined manually based on measurements of the agent’s performance, ensuring that the learning agent receives the necessary signals to learn the correct behavior. Designing a reward function by hand, however, is challenging. Success is hard to formally define and measure in many applications. Beyond that, a sparse signal of success may not be well suited for agent learning – resulting in the need for reward shaping [194], where the reward signal is transformed into one that is more suitable for learning. This often makes the reward signal more susceptible to spurious correlations, however – behaviors that are rewarded because they are usually correlated with the true objective but are not valuable in themselves. This ultimately cumulates in the issue of reward hacking [242], where learning agents exploit reward-specific loopholes to achieve undesired outcomes while still generating high rewards.
@@
RLHF has seen a number of successful applications, advances in methodology, and theoretical insights since the last comparable survey [270]. The applications range from the domain of large language models (LLMs) [197] over image generation [155], continuous control [58] and games [119] to robotics [113]. At the same time, there have been a lot of developments of the methodology since the last comparable survey [270]. Examples of methodological developments include using data augmentation and semi-supervised learning approaches to improve sample complexity [203], using meta-learning to quickly adapt the learned preferences to new tasks [219], fusing multiple feedback types [202], using self-supervised representation learning to increase feedback efficiency [180], actively synthesizing hypothetical behaviors for queries [216], and optimizing queries for ease of answer [30]. Finally, there have been some achievements with regard to theoretical results for the field of RLHF, providing new insights but also new questions for the fundamental mathematical problems underlying the modeling of the learning scenario in RLHF.
In this survey, we, therefore, discuss the current state of affairs with regard to the ongoing research in RLHF, classify the current approaches as well as concisely describe their main characteristics, and give a brief overview of the application areas.
@@ 1.1 Why Human Feedback
In conventional RL, the agent’s objective is defined by a reward function which it aims to maximize [248]. Specifying this reward function can be challenging, particularly in complex domains: What would be a suitable reward function for a robot assisting humans in a household environment or for autonomous vehicles navigating through a busy urban environment? Moreover, even well-defined reward functions can lead to unexpected behaviors due to distributional shifts or over-optimization, raising practical and safety concerns. Learning the agent’s objective from human feedback circumvents reward engineering challenges and fosters robust training, with the reward function dynamically refined and adjusted to distributional shifts as the agent learns.
Feedback vs. Demonstrations
The field of inverse RL aims to infer reward functions from human demonstrations [7]. While this can partially resolve reward engineering challenges, it faces inherent difficulties: (i) It is generally not possible to robustly identify rewards from demonstrations [43], (ii) it is only applicable in scenarios where good demonstrations can be obtained, (iii) it struggles to outperform the demonstrator, and (iv) humans often do not demonstrate the behavior they would prefer a machine to adopt [20]. Interactive feedback, by contrast, can use active queries to differentiate between human preferences and irrelevant noise, is much easier to provide than demonstrations, does not require near-optimal performance from the human evaluators, and elicits preferences on the behavior that a human would prefer from the machine. Interactive feedback can also be used to complement demonstrations, in which case it can be used to shape and refine capabilities learned through initial training, like behavioral cloning, thereby preventing overfitting to demonstrated behavior [2].
@@
Reward engineering in RL presents significant challenges, as accurately specifying reward functions is notoriously difficult [5, 142]. These challenges can be mitigated by utilizing human feedback, which enables training agents for tasks that are hard to define manually and helps avoid safety issues arising from misaligned rewards [242]. Safety issues related to a misalignment between the agent’s and the human’s objectives are studied as the AI alignment problem [85], in particular agent alignment and value alignment [140]. Although the effectiveness of RLHF in resolving these alignment issues is debated [57], it presents a promising approach to enhance alignment [159].
@@
Excessive optimization for poorly specified rewards often leads to unintended behaviors. Agents may exploit simulation flaws for higher rewards [158, 14] or engage in reward hacking [242], where the behavior maximizes the specified reward but deviates from the intended objective. This is evident in cases where agents focus on intermediate rewards without achieving the actual goal [59] or prematurely exit a game to avoid negative rewards [226]. The root of these issues is that the reward function does not properly reflect the actual learning task. While these issues may seem trivial in game-like environments, their implications are far more serious in safety-critical contexts such as healthcare and autonomous driving. In these settings, it is crucial to prevent misaligned reward functions from leading to harmful outcomes, like a care robot causing injury or a self-driving car jeopardizing road safety.
1.2 The Origins of Reinforcement Learning from Human Feedback
@@
Learning behavior from human feedback has long been studied as a subfield of RL, but methods and terminology have evolved over time. Early methods, as discussed in more detail by [141], focused on learning directly from human rewards [121, 144]. This survey, however, focuses on more indirect approaches of inferring the objective from human feedback.
@@
Reinforcement learning from human feedback (RLHF) in its modern guise has its origin in the setting of preference-based reinforcement learning (PbRL) as introduced independently by [3] and [53]. The original idea of preference-based reinforcement learning (PbRL) is to infer the objective from qualitative feedback, such as pairwise preferences between behaviors or between actions given states, instead of quantitiative feedback in the form of numerical rewards. The term RLHF was coined as an alternative later on [9, 200, 197], though initially referring to the same concept of learning behavior from relative feedback.
@@
Disentangling PbRL and RLHF is challenging due to their overlapping use in the literature. For instance, [58] themselves are using the term PbRL, yet are often cited as a seminal reference for RLHF [65, 200]. This indicates the interchangeability of these terms. Practically, RLHF is often associated with reward modeling and deep RL, while PbRL is often linked to direct policy optimization in traditional RL settings. This is underlined by [125], who characterize PbRL as limited to direct policy learning from preferences. This is in contrast with other sources, however, who include reward learning within the scope of RLHF [58, 270].
Despite the overlapping and sometimes conflicting usage, RLHF is increasingly viewed as a generalization of PbRL. While both involve human feedback to define RL objectives, PbRL primarily focuses on relative feedback, such as binary comparisons and rankings. RLHF not only includes these aspects but also extends to a wider range of feedback types [181]. Table 1 gives an exemplary overview of our interpretation of these terms.
Another concept, semi-supervised reinforcement learning (SSRL), introduced by [56] and discussed by [5], refers to an RL setting where an agent receives feedback on a subset of its experiences. The initial discussions of SSRL focused on absolute feedback on subsets of the agent’s experiences, making the concept complementary to PbRL. In contrast to PbRL and RLHF, the term SSRL seems to be used less in the recent literature.
@@ 1.3 Scope of the Survey
We focus on approaches that learn a reward model from human feedback and then use this model to train a policy. Although it is possible to directly optimize a policy from human feedback [270], thereby performing RLHF without reward learning, this approach has been practiced only rarely so far. The decomposition into reward learning and policy training offers many conceptual and practical benefits. Among those benefits are the direct applicability of supervised learning techniques for the reward model and the possibility of evaluating the reward model in isolation. In addition to that, the decomposition naturally leads to a form of semi-supervised learning, enabling the agent to use labeled episodes for reward model training while leveraging unlabelled episodes to refine its behavior and explore the environment. Note, however, that there is a recent trend in the direction of direct policy learning in the domain of language model fine-tuning (see Section 5.2.3).
@@
In addition to these criteria, we mainly focus on works published after 2017 since earlier works are surveyed by [270]. Nevertheless, some works from this period are revisited from time to time in order to elaborate on certain concepts that are still state of the art or have significantly shaped it. Exceptions are made if the used methods are of interest for RLHF approaches.
@@ 1.4.1 Human-in-the-Loop RL
[275]
HITL machine learning.
[190]
RL with human advice.
[164]
Social feedback.
[207]
Human brain signals.
[8]
Interactive RL for HCI.
[199]
Imitation learning.
[7]
Inverse RL.
[27]
Assisted RL.
[174]
Language-informed RL.
[290]
Human guidance.
[126]
AI Alignment.
[170]
RLHF for LLMs.
Ours
RLHF.
Human participation in machine learning, particularly in guiding machine learners, is a much-studied scenario. This field, commonly referred to as human-in-the-loop (HITL) machine learning, can be further divided into subfields based on various criteria, e.g., the ones detailed in Section 1.3. Prior surveys of these subfields are compiled in Table 2 and briefly summarized in the following.
@@
[275] survey HITL machine learning in general. They also cover some applications of RLHF (for LLMs in particular) but do not give a detailed overview. Similarly broad in scope, [190] study the setting of RL with human advice, which they define as ‘teaching signals that can be communicated by the teacher to the learning system without executing the task.’ While this setting subsumes RLHF, the broad generality limits the depth in which their survey can cover RLHF approaches.
@@
[8] survey interactive RL from an human-computer interaction (HCI) viewpoint, exploring various ways humans can influence RL agents, with a particular focus on reward definition based on human feedback, without a predefined environmental reward function. Due to the breadth of their survey, they do not cover many works in this area. The survey by [164] centers on interactive RL using human social cues, like gestures and spoken language, but does not cover the reward modeling aspect. Similarly, the study by [207] examines RL with direct feedback from human brain signals, such as through brain-computer interfaces, also not focusing on reward modeling.
@@
Learning from demonstrations, in the form of behavior cloning [199] and inverse RL [7], shares the goal of RLHF to learn behavior from human input. In contrast to RLHF, however, it requires demonstrations of the desired behavior instead of feedback, and these demonstrations are usually not provided interactively and online. This limits their applications and also their final performance due to the availability of near-optimal demonstrations. Nonetheless, imitation and demonstration can be a useful component of an RLHF system but are not the main focus of this survey. However, we will discuss the intersection between these fields in some parts whenever necessary.
@@
[27] review the field of assisted RL, where an agent may receive external information (for example, from a human) that aids it in action selection. While updates to the reward function are one of the possible effects of advice in this setting (in addition to action selection or modifications of the agent’s internal state), it is usually assumed that an initial reward function is given and the extent of the updates is limited to reward shaping or supplementary reward signals. In contrast to RLHF, the external information does not define the task but only helps the agent in achieving it. Closely related to this, [174] survey RL assisted by natural language. In addition to this assistance setting, they also discuss approaches that infer a language-conditioned reward function. However, they discuss this setting rather briefly and use techniques from inverse RL and not RLHF.
@@
In their survey on human guidance, [290] delve into various aspects related to RLHF. Although they touch on aspects such as reward learning, it is not the primary emphasis of their work. Instead, their main focus lies on exploring more immediate approaches that do not involve the learning of a reward model.
@@
[126] provide a general overview of AI alignment, i.e., the challenge of aligning the objectives of an intelligent system with those of its human operators. This survey covers RLHF in some detail. As AI alignment is a very broad field, however, the article nevertheless does not go into as much depth on the topic of RLHF as we do here.
@@
[170] give an overview of current applications of RLHF methods for LLMs such as ChatGPT and GPT-4. Even though it currently enjoys a lot of attention, it is only a specific application area for RLHF. Our survey adopts a broader perspective, examining the diverse applications and impact of RLHF encompassing application areas beyond LLMs.

### v2

# algorithmic

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a variant of reinforcement learning (RL) that learns from human feedback instead of relying on an engineered reward function. Building on prior work on the related setting of preference-based reinforcement learning (PbRL), it stands at the intersection of artificial intelligence and human-computer interaction. This positioning offers a promising avenue to enhance the performance and adaptability of intelligent systems while also improving the alignment of their objectives with human values. The training of large language models (LLMs) has impressively demonstrated this potential in recent years, where RLHF played a decisive role in directing the model’s capabilities toward human objectives. This article provides a comprehensive overview of the fundamentals of RLHF, exploring the intricate dynamics between RL agents and human input. While recent focus has been on RLHF for LLMs, our survey adopts a broader perspective, examining the diverse applications and wide-ranging impact of the technique. We delve into the core principles that underpin RLHF, shedding light on the symbiotic relationship between algorithms and human feedback, and discuss the main research trends in the field. By synthesizing the current landscape of RLHF research, this article aims to provide researchers as well as practitioners with a comprehensive understanding of this rapidly growing field of research.
@@ 1 Introduction
In reinforcement learning (RL), an agent traditionally navigates through an environment and attempts to make optimal decisions (i.e., action choices) through a process of trial and error. Whether a decision is optimal or not is determined solely by reward signals. These signals have to be defined by a system designer based on measurements of the agent’s performance, ensuring that the learning agent receives the necessary feedback to learn the correct behavior. Designing a reward function, however, is challenging. Indeed, success is hard to formally define and measure in many applications. Beyond that, a sparse signal of success may not be well suited for agent learning – resulting in the need for reward shaping (Ng et al., 1999), where the reward signal is transformed into one that is more suitable for learning. This often makes the reward signal more susceptible to spurious correlations, however – behaviors that are rewarded because they are usually correlated with the true objective but are not valuable in themselves. This ultimately cumulates in the issue of reward hacking (Skalse et al., 2022), where learning agents exploit reward-specific loopholes to achieve undesired outcomes while still generating high rewards.
@@
RLHF has seen a number of successful applications, advances in methodology, and theoretical insights since the last comparable survey (Wirth et al., 2017). The applications span various domains, including large language model (LLM) fine-tuning (OpenAI, 2022), image generation (Lee et al., 2023), continuous control (Christiano et al., 2017), games (Ibarz et al., 2018), and robotics (Hejna & Sadigh, 2022). At the same time, there have been a lot of developments in the methodology since the last comparable survey (Wirth et al., 2017). Examples of methodological developments include fusing multiple feedback types to leverage their relative strengths (see Section 3.5), enhancing query efficiency through active learning and active query synthesis (see Section 4.1.1), incorporating psychological insights to improve the quality of human feedback (see Section 4.2.1), using techniques such as meta-learning to quickly adapt learned preferences to new tasks using prior data (see Section 5.5.1), and using available preference data more efficiently through approaches such as data augmentation and semi-supervised learning (see Section 5.5.2). Finally, there have been some achievements with regard to theoretical results for the field of RLHF (Section 7), providing new insights but also new questions for the fundamental mathematical problems underlying the modeling of the learning scenario in RLHF.
In this survey, we, therefore, discuss the current state of affairs with regard to the ongoing research in RLHF, classify the current approaches as well as concisely describe their main characteristics, and give a brief overview of the application areas. In the remainder of this section, we will start by discussing the motivation (Section 1.1) and origins (Section 1.2) of RLHF as well as the scope of this survey (Section 1.3) and conclude by outlining the contents of the following sections (Section 1.5).
@@ 1.1 Why Human Feedback
In conventional RL, the agent’s objective is defined by a reward function that it aims to maximize (Sutton & Barto, 2018). Specifying this reward function can be challenging, particularly in complex domains: What would be a suitable reward function for a robot assisting humans in a household environment or for autonomous vehicles navigating through a busy urban environment? Moreover, even reward functions that initially seem well-defined can lead to unexpected behaviors due to distributional shifts or over-optimization, raising practical and safety concerns. Learning the agent’s objective from human feedback circumvents reward engineering challenges and fosters robust training, with the reward function dynamically refined and adjusted to distributional shifts as the agent learns.
Interactive Feedback vs. Demonstrations
The field of inverse RL aims to infer reward functions from human demonstrations (Arora & Doshi, 2021). While this can partially resolve reward engineering challenges, it faces inherent difficulties: (i) It is generally not possible to robustly identify rewards from demonstrations (Cao et al., 2021a; Mindermann & Armstrong, 2018), (ii) it is only applicable in scenarios where good demonstrations can be obtained, (iii) it struggles to outperform the demonstrator, and (iv) humans often do not demonstrate the behavior they would prefer a machine to adopt (Basu et al., 2017). Interactive feedback, by contrast, can use active queries to differentiate between human preferences and irrelevant noise, is much easier to provide than demonstrations, does not require near-optimal performance from the human evaluators, and elicits preferences on the behavior that a human would prefer from the machine. Interactive feedback can also be used to complement demonstrations, in which case it can be used to shape and refine capabilities learned through initial training, like behavioral cloning, thereby preventing overfitting to demonstrated behavior (Abramson et al., 2022).
@@
Reward engineering in RL presents significant challenges, as accurately specifying reward functions is notoriously difficult (Amodei et al., 2016; Knox et al., 2023). These challenges can be mitigated by utilizing human feedback, which enables training agents for tasks that are hard to define manually and helps avoid safety issues arising from misaligned rewards (Skalse et al., 2022). Safety issues related to a misalignment between the agent’s and the human’s objectives are studied as the AI alignment problem (Gabriel, 2020), in particular agent alignment and value alignment (Kirchner et al., 2022).
@@
Excessive optimization for poorly specified rewards often leads to unintended behaviors. Examples of such behaviors include exploiting flaws in the environment simulation for higher rewards (Lehman et al., 2020; Baker et al., 2020) or engaging in more general reward hacking (Skalse et al., 2022), where the behavior maximizes the specified reward but deviates from the intended objective. This is evident in cases where agents focus on intermediate rewards without achieving the actual goal (Clark & Amodei, 2016) or prematurely exit a game to avoid negative rewards (Saunders et al., 2018). The root of these issues is that the reward function does not properly reflect the actual learning task. While these issues may seem trivial in game-like environments, their implications are far more serious in safety-critical contexts such as healthcare and autonomous driving. In these settings, it is crucial to prevent misaligned reward functions from leading to harmful outcomes, like a care robot causing injury or a self-driving car jeopardizing road safety.
RLHF presents a promising approach to enhance alignment (Leike et al., 2018) by enabling agents to learn from human feedback, which is often more closely aligned with the true objective than manually specified rewards. Nonetheless, the effectiveness of RLHF in resolving these alignment issues is debated (Christiano, 2023). Examples of possible pitfalls raised in this debate are that the agent may be incentivized to manipulate the human teacher to provide feedback that is easier to optimize (Armstrong et al., 2020; Carroll et al., 2023) or that the agent may learn to exploit errors in human judgement (Ngo et al., 2024). We refer the interested reader to the survey by Casper et al. (2023) for a more detailed discussion of these issues. Despite this debate, RLHF represents an important early step towards aligning agents with human values and serves as a foundation to build on to further improve agent alignment.
1.2 The Origins of RLHF
@@
Learning behavior from human feedback has long been studied as a subfield of RL, but methods and terminology have evolved over time. Early methods focused on learning directly from human rewards (Knox, 2012; Isbell et al., 2001; Knox & Stone, 2009), action advice (Maclin et al., 2005), or action critique (Judah et al., 2010). Notable approaches in this area include TAMER (Knox & Stone, 2009; Warnell et al., 2018), which interprets human feedback as samples of the optimal action-value function, and the later COACH (MacGlashan et al., 2017; Arumugam et al., 2019), which interprets human feedback in a policy-dependent way, i.e., as samples of the advantage function. This survey, however, focuses on more indirect approaches to inferring the objective from human feedback.
@@
Reinforcement learning from human feedback (RLHF) in its modern guise has its origin in the setting of preference-based reinforcement learning (PbRL) as introduced independently by Akrour et al. (2011) and Cheng et al. (2011). The original idea of preference-based reinforcement learning (PbRL) is to infer the objective from qualitative feedback, such as pairwise preferences between behaviors or between actions given states, instead of quantitative feedback in the form of numerical rewards. The term RLHF was coined as an alternative later on (Askell et al., 2021; Ouyang et al., 2022; OpenAI, 2022), though initially referring to the same concept of learning behavior from relative feedback.
@@
Disentangling PbRL and RLHF is challenging due to their overlapping use in the literature. For instance, Christiano et al. (2017) themselves are using the term PbRL, yet are often cited as a seminal reference for RLHF (Daniels-Koch & Freedman, 2022; Ouyang et al., 2022). This indicates the interchangeability of these terms. Practically, RLHF is often associated with reward modeling and deep RL, while PbRL is often linked to direct policy optimization in traditional RL settings. This is underlined by Jeon et al. (2020), who characterize PbRL as limited to direct policy learning from preferences. This is in contrast with other sources, however, who include reward learning within the scope of PbRL (Christiano et al., 2017; Wirth et al., 2017).
Despite the overlapping and sometimes conflicting usage, RLHF is increasingly viewed as a generalization of PbRL. While both involve human feedback to define RL objectives, PbRL primarily focuses on relative feedback, such as binary comparisons and rankings. RLHF not only includes these aspects but also extends to a wider range of feedback types (Metz et al., 2023; Yuan et al., 2024). Table 1 gives an exemplary overview of our interpretation of these terms.
Another concept, semi-supervised reinforcement learning (SSRL), introduced by Christiano (2016) and discussed by Amodei et al. (2016), refers to an RL setting where an agent receives feedback on a subset of its experiences. The initial discussions of SSRL focused on absolute feedback on subsets of the agent’s experiences, making the concept complementary to PbRL. In contrast to PbRL and RLHF, the term SSRL seems to be used less in the recent literature.
@@ 1.3 Scope of the Survey
We focus on approaches that learn a reward model from human feedback and then use this model to train a policy. Although it is possible to directly optimize a policy from human feedback (Wirth et al., 2017), thereby performing RLHF without reward learning, this approach was almost not practiced for a long time and has only recently gained renewed interest especially in the domain of language model fine-tuning (see Section 6.3). The decomposition into reward learning and policy training offers many conceptual and practical benefits. Among those benefits are the direct applicability of supervised learning techniques for the reward model and the possibility of evaluating the reward model in isolation. In addition to that, the decomposition naturally leads to a form of semi-supervised learning, enabling the agent to use labeled episodes for reward model training while leveraging unlabelled episodes to refine its behavior and explore the environment.
@@
In addition to these criteria, we mainly focus on works published after 2017 since earlier works are surveyed by Wirth et al. (2017). Nevertheless, some works from this period are revisited from time to time in order to elaborate on certain concepts that are still state of the art or have significantly shaped it. Note that while there has been a recent forus on RLHF for LLMs, LLMs are not the primary focus of this work. Instead, we cover RLHF in a broader context, focusing on control applications as well as discussing the implications of RLHF for fine-tuning generative models such as LLMs.
@@ 1.4.1 Human-in-the-Loop RL
Wu et al. (2022) Human-in-the-loop ML
Retzlaff et al. (2024) Human-in-the-loop RL (✓) (✓) (✓)
Najar & Chetouani (2021) RL with human advice (✓) (✓) (✓)
Lin et al. (2020a) Social feedback (✓)
Poole & Lee (2024) RL from brain signals
Cruz & Igarashi (2020) Interactive RL for HCI
Osa et al. (2018) Imitation learning
Arora & Doshi (2021) Inverse RL
Bignold et al. (2021) Assisted RL (✓)
Luketina et al. (2019) Language-informed RL
Zhang et al. (2021) Human guidance
Ji et al. (2023b) AI Alignment
Liu et al. (2023c) LLM applications
Ours RLHF
Human participation in machine learning (ML), particularly in guiding machine learners, is a much-studied scenario. This field, commonly referred to as human-in-the-loop ML, can be further divided into subfields based on various criteria, e.g., the ones detailed in Section 1.3. Prior surveys of these subfields are compiled in Table 2 and briefly summarized in the following.
@@
Learning from human feedback falls into the domain of human-in-the-loop ML. Wu et al. (2022) survey human-in-the-loop ML in general. They also cover some applications of RLHF (for LLMs in particular) but do not give a detailed overview. Retzlaff et al. (2024) provide a similar overview over human-in-the-loop RL in particular, focusing on human involvement in RL on a more abstract level than our work and not covering RLHF in detail. Similarly broad in scope, Najar & Chetouani (2021) study the setting of RL with human advice, which they define as ‘teaching signals that can be communicated by the teacher to the learning system without executing the task.’ While this setting subsumes RLHF, the broad generality limits the depth in which their survey can cover RLHF approaches.
@@
Cruz & Igarashi (2020) survey interactive RL from an human-computer interaction (HCI) viewpoint, exploring various ways humans can influence RL agents, with a particular focus on reward definition based on human feedback, without a predefined environmental reward function. Due to the breadth of their survey, they do not cover many works in this area. The survey by Lin et al. (2020a) centers on interactive RL using human social cues, like gestures and spoken language, but does not cover the reward modeling aspect. Similarly, the study by Poole & Lee (2024) examines RL with direct feedback from human brain signals, such as through brain-computer interfaces, also not focusing on reward modeling.
@@
Learning from demonstrations, in the form of behavior cloning (Osa et al., 2018) and inverse RL (Arora & Doshi, 2021), shares the goal of RLHF to learn behavior from human input. In contrast to RLHF, however, it requires demonstrations of the desired behavior instead of feedback, and these demonstrations are usually not provided interactively and online. This limits their applications and also their final performance due to the availability of near-optimal demonstrations. Nonetheless, imitation and demonstration can be a useful component of an RLHF system but are not the main focus of this survey. However, we will discuss the intersection between these fields in some parts whenever necessary.
@@
Bignold et al. (2021) review the field of assisted RL, where an agent may receive external information (for example, from a human) that aids it in action selection. While updates to the reward function are one of the possible effects of advice in this setting (in addition to action selection or modifications of the agent’s internal state), it is usually assumed that an initial reward function is given and the extent of the updates is limited to reward shaping or supplementary reward signals. In contrast to RLHF, the external information does not define the task but only helps the agent in achieving it. Closely related to this, Luketina et al. (2019) survey RL assisted by natural language. In addition to this assistance setting, they also discuss approaches that infer a language-conditioned reward function. However, they discuss this setting rather briefly and use techniques from inverse RL and not RLHF.
@@
In their survey on human guidance, Zhang et al. (2021) delve into various aspects related to RLHF. Although they touch on aspects such as reward learning, it is not the primary emphasis of their work. Instead, their main focus lies on exploring more immediate approaches that do not involve the learning of a reward model.
@@
Ji et al. (2023b) provide a general overview of AI alignment, i.e., the challenge of aligning the objectives of an intelligent system with those of its human operators. This survey covers RLHF in some detail. As AI alignment is a very broad field, however, the article nevertheless does not go into as much depth on the topic of RLHF as we do here.
@@
Liu et al. (2023c) give an overview of current applications of RLHF methods for LLMs such as ChatGPT and GPT-4. Even though it currently enjoys a lot of attention, it is only one specific application area for RLHF. Our survey adopts a broader perspective, examining the diverse applications and impact of RLHF encompassing application areas beyond LLMs.